An ice freezer behind a restaurant has a freon leak, releasing 36.23 g of C2H2F3Cl into the air every week. If the leak is not fixed, how many kilograms of fluorine will be released into the air over 6 months? Assume there are 4 weeks in a month.
1
Expert's answer
2019-10-24T08:00:27-0400

Total number of weeks"=6\times 4=24"

Mass leak rate of freon"=36.23 Kg\/week"

Mass leak rate of flourine"= \frac{Flourine \ mass \ in \ Freon}{ M.M.\ Of \ Freon}\times leak\ rate=" "\frac{19\times 3}{118.5}\times 36.23=17.427\ gm/week"

Total flourine leaked in 6 months"=24\times17.427=418.248\ gm=0.418 Kg"